# Drivers of cooperation in social dilemmas on higher-order networks

### v1

@@ I Introduction
Usually, social dilemmas comprise of two interacting individuals, each of whom earns a payoff based on their strategies, typically denoted as cooperation (C) and defection (D). Combining game theory with the Darwinian principle of the survival of the fittest, evolutionary game theory extends this paradigm to consider repeated interactions in a population. The fraction of individuals adopting each strategy changes over time based on how fit the strategy is in the population, where the fitness is determined by the average payoff earned by players adopting that particular strategy Axelrod and Hamilton (1981); Nowak and May (1992); Hofbauer and Sigmund (1998); Traulsen and Glynatsi (2023). Over the years extensive numerical and theoretical work for various games and social dilemma situations have identified several different mechanisms contributing to the success of cooperative behaviour. These mechanisms can be broadly classified into reciprocity based (direct Trivers (1971); Axelrod and Hamilton (1981), indirect Nowak and Sigmund (1998); Hilbe et al. (2018), or network Lieberman et al. (2005); Szabó and Fáth (2007)) or selection based (kin Hamilton (1964); Taylor (1992) or group Wilson (1975); Traulsen and Nowak (2006)) as pointed out in Nowak (2006b).
@@
Repeated interactions with the same set of individuals (or neighbours) is at the heart of network reciprocity which has gathered substantial attention in recent years as a computational and mathematical framework able to explain cooperation in socio-structural interaction patterns Hofbauer and Sigmund (1998); Santos and Pacheco (2005); Lieberman et al. (2005); Nowak (2006b); Santos et al. (2006); Szabó and Fáth (2007); Gómez-Gardeñes et al. (2007). In particular, the effect of different interaction structures has been investigated, from lattices Nowak and May (1992); Szabó and Fáth (2007); Kumar et al. (2020), to scale-free networks Santos and Pacheco (2005); Santos et al. (2008), and multi-layer networks Wang et al. (2015); Battiston et al. (2017); Su et al. (2022) enhancing our understanding of how pro-sociality evolves in large-scale structured populations with realistic features. A crucial limitation of networks is that they encode ties with links, and are hence able to properly represent pairwise interactions only. By contrast, group sociality is a ubiquitous feature of human interaction, highlighting the need for frameworks that can accommodate multi-player interactions within structured populations Broom et al. (1997); Tarnita et al. (2009); Gokhale and Traulsen (2010); Broom et al. (2019).
@@
Hypergraphs (or higher-order networks) are mathematical objects that encode group interactions through hyperedges, describing interactions among an arbitrary number of individuals Battiston et al. (2020, 2021). A variety of dynamical processes such as spreading of behaviours Iacopini et al. (2019), synchronization of coupled oscillators Skardal and Arenas (2020); Gambuzza et al. (2021); Zhang et al. (2023), and very recently evolutionary games Civilini et al. (2021); Alvarez-Rodriguez et al. (2021); Guo et al. (2021), have been shown to exhibit fundamentally different behaviour when considered on hypergraphs. In particular, a higher-order prisoner’s dilemma game was shown to exhibit an explosive transition to cooperation at a critical fraction of higher-order interactions Civilini et al. (2024). Similar works on higher-order networks for other types of games such as public goods game Wang et al. (2024) and sender-receiver game Kumar et al. (2021) have shown analogous non-trivial behaviour with respect to the topological properties of the hypergraphs.
@@ II Model
A population of $N$ players is represented as a hypergraph $\mathcal{H}(\mathcal{V},\mathcal{E})$, where the players are the nodes of the hypergraph $\mathcal{V}$, such that $|\mathcal{V}|=N$ and the set of hyperedges $\mathcal{E}$ represents the set of games played in the population. Hyperedges are a generalization of the network’s edges to group interactions, a hyperedge representing a group of an arbitrary number of interacting nodes. In particular, we focus on the case where the hypergraph $\mathcal{H}$ consists of only $2-$hyperedges and $3-$hyperedges, respectively denoted by $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}$. A $2$-player game is associated to each $2-$hyperedge, while a $3$-player game to each $3-$hyperedge. Henceforth, we will use superscripts /// and $\Delta$ to indicate pairwise and higher-order quantities respectively, while reserving the subscripts for node labels. Each player participates in all the hyperedges to which it belongs. Since each player is part of multiple hyperedges, they engage in a separate game for each one. Thus, the total number of games a player $i$ takes part in equals their hyperdegree $k_{i}$. We define $k_{i}^{/}$ and $k_{i}^{\Delta}$ as the number of 2-player and 3-player games player $i$ is involved in, respectively, satisfying $k_{i}^{/}+k_{i}^{\Delta}=k_{i}$.
@@ II.1 Payoffs
where the first entry in each cell denotes the payoff earned by the row player, while the second entry is the payoff for the column player. $R^{/}$ denotes the reward for mutual cooperation, while $P^{/}$ denotes the penalty for mutual defection. A cooperator earns a payoff $S^{/}$ (namely sucker’s payoff) against a defector, while the defector earns a payoff $T^{/}$ (temptation to defect) when facing a cooperator. The relative order of these payoffs determines the type of 2-player social dilemmas. In particular for our analysis, we will study prisoner’s dilemma games, defined by the ordering $T^{/}>R^{/}>P^{/}>S^{/}$. To reduce the number of parameters in the game, we fix $R^{/}=1$ and $P^{/}=0$ to define the limit of the parameters as is usually done in the literature Szabó and Fáth (2007); Guo et al. (2021); Wang et al. (2023); Civilini et al. (2024).
@@
As mentioned in the introduction, a social dilemma is a type of collective action problem where there is a tension between personal benefit and collective good. For instance, a $2$-player Prisoner’s dilemma represents a social dilemma since, given the payoff ordering $T^{/}>R^{/}>P^{/}>S^{/}$ defining the game, the rational choice (or Nash equilibrium) for a player is to choose defection, even though it would be more beneficial (i.e., it would bring a higher payoff) to cooperate for both players. However, generalizing the concept of social dilemma to multi-player games is not trivial. Over the years different definitions Kerr et al. (2004); Peña et al. (2016) of social dilemmas in multi-player game representation have been proposed. We will adopt the definition proposed by Ref. Broom et al. (2019), according to which the payoffs of a $3$-player game have to satisfy the following conditions to qualify as a social dilemma:
@@
However, in the higher-order game, two additional payoffs, $G^{\Delta}$ and $W^{\Delta}$, are introduced. If $G^{\Delta}<W^{\Delta}$, all the above conditions (a-d) hold, classifying the game as a strong social dilemma. Conversely, when $G^{\Delta}>W^{\Delta}$, condition d is not met, leading to what is known as a relaxed social dilemma. In general, the quantity $\alpha=W^{\Delta}-G^{\Delta}$ can be interpreted as a measure of the strength of the higher-order Prisoner’s Dilemma (PD). Notably, the strong social dilemma has only full defection ($\{D,D,D\}$) as a Nash equilibrium, whereas in the relaxed social dilemma, besides full defection, the strategy profile $\{C,C,D\}$ also forms a Nash equilibrium Gokhale and Traulsen (2010); Hilbe et al. (2014); Civilini et al. (2024).
@@
It is crucial that payoffs for hyperedges of different sizes remain comparable so that larger groups do not give an undue advantage (or disadvantage) simply because of their group interaction Capraro and Barcelo (2015); Peña and Nöldeke (2018); Pereda et al. (2019); Wang et al. (2024). For instance, consider a situation where a player cooperates with two cooperating players in two dyadic interactions earning a total payoff of $2R^{/}$. If instead the player cooperated with the same two cooperating players but in a 3-player interaction, it will earn a payoff of $R^{\Delta}$. Since the evolutionary dynamics typically consists of aggregating payoffs to determine the fitness of strategies, if $2R^{/}\neq R^{\Delta}$, the total payoff earned by a player by interacting with the same players but with different types of interaction (pairwise or higher-order) will be different. Even though larger groups can have synergistic effects Capraro and Barcelo (2015); Peña and Nöldeke (2018); Pereda et al. (2019); Lee et al. (2023) making cooperation naturally more or less advantageous, we want to focus on emergent collective behaviour in absence of these phenomena. Inspired by this line of thinking, one can imagine that from a player’s perspective, a single 3-player interaction is equivalent to two pairwise interactions. Here, we introduce a new payoff constraint for the pairwise and higher-order payoffs, $R^{\Delta}\sim R^{/}+R^{/}=2R^{/}$ and similarly for $S^{\Delta}=2S^{/}$, $T^{\Delta}=2T^{/}$, and $P^{\Delta}=2P^{/}$. We tune the values of $G^{\Delta}$ and $W^{\Delta}$ to explore the landscape of higher-order PD based on its strength mentioned above.
@@ II.2 Topological overlap in higher-order networks
We consider a random-regular hypergraph such that a player $i$ has $k^{/}$ 2-hyperedges (or pairwise neighbours) and further participates in $k^{\Delta}$ 3-hyperedges (or triangles). For our simulations, we fix $k^{/}=4$ and $k^{\Delta}=2$ so that each player interacts with maximum 8 (=$k^{/}+2k^{\Delta})$ other unique players. We can represent any form of higher-order interactions as a multilayer hypergraph, where each layer consists of interactions of a specific size/order, as illustrated in fig. (1). The advantage of such a representation is that we can adapt well-established measures from the theory of multiplex networks to characterize the structure of our system. We are interested in exploring how topological similarity between the pairwise and higher-order interactions impacts the emergence of cooperation in structured populations Battiston et al. (2014); Lee et al. (2021); Anwar and Ghosh (2022); Zhang et al. (2023); Malizia et al. (2025); Presigny et al. (2024).
We denote $\mathcal{E}^{\Delta}_{\text{proj}}$ as the set of projection of 3-hyperedges unto their corresponding $2$-hyperedges. In other words, if $e_{lmn}\in\mathcal{E}^{\Delta}$, then $e_{lm},e_{mn},e_{ln}\in\mathcal{E}^{\Delta}_{\text{proj}}$. We then define the topological overlap $\omega$ between $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}$ as the size of the intersection between the sets of edges $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}_{\text{proj}}$ normalized by the size of set of edges in the pairwise network Battiston et al. (2014, 2017); Krishnagopal and Bianconi (2023),
@@ II.3 Evolutionary dynamics
To account for the dynamical coupling between interaction orders, we assume that the focal player can imitate the strategies of the model player associated with the two layers. To illustrate with an example consider the update of the pairwise strategy of a focal player $f$ denoted by $s_{f}^{/}$ (fig. (1)). We now assume that the player can copy the pairwise strategy of the model player $m$ denoted by $s_{m}^{/}$. However, here we propose that the dynamical coupling between the two layers also enables the focal player to copy the higher-order strategy of the model player denoted by $s_{m}^{\Delta}$. We denote the probability of imitating the strategy of the model player from the other layer as $p_{\text{switch}}$, controlling the dynamical coupling between the two layers and allowing for inter-order imitation. The profile of all transition probabilities at a given time step can be summarised as follows,
@@
We obtain one full Monte-Carlo Step (MCS) by repeating the above procedure $2\cdot N$ times so that each player gets the opportunity to update both the pairwise and higher-order strategies once on average. The structural and dynamical ingredients of our model are fully visualized in fig. (1).
We characterize the pro-social behaviour in our system using a variety of descriptors. We introduce $\rho_{0}^{/}$ and $\rho_{0}^{\Delta}$ as the initial density of cooperators in the pairwise and higher-order layers respectively. Correspondingly, $\rho_{0}=\frac{1}{2}[\rho_{0}^{/}+\rho_{0}^{\Delta}]$ denotes the initial mass of cooperators in the system. We always consider $\rho_{0}^{/}=\rho_{0}^{\Delta}=\rho_{0}=0.5$ unless stated otherwise. To quantify the stationary state properties of the system, i.e. a state where the observables of the system become time independent, we compute $\rho^{/}$ as the fraction of nodes cooperating in the pairwise interactions and $\rho^{\Delta}$ as the fraction of nodes cooperating in the higher-order interactions. We denote the overall cooperation level in the system as Wang et al. (2015); Battiston et al. (2017),
@@ III.1 Inter-order dynamical coupling mediates cooperation
We first investigate the effect of the dynamical coupling between the different orders of interactions in our model by tuning $p_{\text{switch}}$. Fig. (2) depicts the stationary state cooperation levels $\langle\rho\rangle$ as a function of $p_{\text{switch}}$ for various values of social dilemma strength $\alpha$ and structural overlap $\omega$ in multi-dimensional strategy systems. To highlight the advantages of our model, we also plot in coloured dotted lines, the cooperation levels in a similar system except with a single/scalar strategy assigned to each node instead of a vector of strategies. In other words, we showcase the difference between the presence and absence of group-size-dependent strategies. Note that in the case where each node has only one associated strategy, the notion of $p_{\text{switch}}$ is meaningless. First, we observe that depending on the value of $\alpha$, the uni-dimensional scalar-strategy system shows a transition from full defection to high levels of cooperation Civilini et al. (2024). Second, we notice that for both values of overlap $\omega=0.0$ and $\omega=1.0$, for $p_{\text{switch}}=0$, the system displays 50% cooperative agents. We can understand this behaviour by noting that the optimum strategy for the pairwise PD game based on our chosen payoff values is pure defection, while for the higher-order game, since $\alpha<0$, one of the Nash equilibria is to cooperate (see Sec. (II.1)) Gokhale and Traulsen (2010); Civilini et al. (2024).

### v2

@@ I Introduction
Usually, social dilemmas comprise of two interacting individuals, each of whom earns a payoff based on their strategies, typically denoted as cooperation (C) and defection (D). Combining game theory with the Darwinian principle of the survival of the fittest, evolutionary game theory extends this paradigm to consider repeated interactions in a population. The fraction of individuals adopting each strategy changes over time based on how fit the strategy is in the population, where the fitness is determined by the average payoff earned by players adopting that particular strategy Axelrod and Hamilton (1981); Nowak and May (1992); Hofbauer and Sigmund (1998); Traulsen and Glynatsi (2023). Over the years, extensive numerical and theoretical work for various games and social dilemma situations have identified several different mechanisms contributing to the success of cooperative behaviour. These mechanisms can be broadly classified into reciprocity based (direct Trivers (1971); Axelrod and Hamilton (1981), indirect Nowak and Sigmund (1998); Hilbe et al. (2018), or network Lieberman et al. (2005); Szabó and Fáth (2007)) or selection based (kin Hamilton (1964); Taylor (1992) or group Wilson (1975); Traulsen and Nowak (2006)) as pointed out in Nowak (2006b).
@@
Repeated interactions with the same set of individuals (or neighbours) is at the heart of network reciprocity which has gathered substantial attention in recent years as a computational and mathematical framework able to explain cooperation in socio-structural interaction patterns Hofbauer and Sigmund (1998); Santos and Pacheco (2005); Lieberman et al. (2005); Nowak (2006b); Santos et al. (2006); Szabó and Fáth (2007); Gómez-Gardeñes et al. (2007). In particular, the effect of different interaction structures has been investigated, from lattices Nowak and May (1992); Szabó and Fáth (2007); Kumar et al. (2020), to scale-free networks Santos and Pacheco (2005); Santos et al. (2008), and multi-layer networks Wang et al. (2015); Battiston et al. (2017); Su et al. (2022) enhancing our understanding of how pro-sociality evolves in large-scale structured populations with realistic features. A crucial limitation of networks is that they encode ties with links, and are hence unable to properly represent non-pairwise interactions. By contrast, group sociality is a ubiquitous feature of human interaction, highlighting the need for frameworks that can accommodate multi-player interactions within structured populations Broom et al. (1997); Tarnita et al. (2009); Gokhale and Traulsen (2010); Broom et al. (2019).
@@
Hypergraphs (or higher-order networks) are mathematical objects that encode group interactions through hyperedges, describing interactions among an arbitrary number of individuals Battiston et al. (2020, 2021). A variety of dynamical processes such as spreading of behaviours Iacopini et al. (2019), synchronization of coupled oscillators Skardal and Arenas (2020); Gambuzza et al. (2021); Zhang et al. (2023), and very recently evolutionary games Civilini et al. (2021); Alvarez-Rodriguez et al. (2021); Guo et al. (2021); Xu et al. (2024), have been shown to exhibit fundamentally different behaviour when considered on hypergraphs. In particular, a higher-order prisoner’s dilemma game was shown to exhibit an explosive transition to cooperation at a critical fraction of higher-order interactions Civilini et al. (2024). Similar works on higher-order networks for other types of games such as public goods game Wang et al. (2024a) and sender-receiver game Kumar et al. (2021) have shown analogous non-trivial behaviour with respect to the topological properties of the hypergraphs.
@@ II Model
A population of $N$ players is represented as a hypergraph $\mathcal{H}(\mathcal{V},\mathcal{E})$, where the players are the nodes of the hypergraph $\mathcal{V}$, such that $|\mathcal{V}|=N$ and the set of hyperedges $\mathcal{E}$ represents the set of games played in the population. Hyperedges are a generalization of the network’s edges to group interactions, a hyperedge representing a group of an arbitrary number of interacting nodes. In particular, we focus on the case where the hypergraph $\mathcal{H}$ consists of only $2-$hyperedges and $3-$hyperedges, respectively denoted by $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}$. A $2$-player game is associated to each $2-$hyperedge, while a $3$-player game to each $3-$hyperedge. Henceforth, we will use superscripts /// and $\Delta$ to indicate pairwise and higher-order quantities respectively, while reserving the subscripts for node labels. Each player participates in all the hyperedges to which it belongs. Since every player is part of multiple hyperedges, they engage in a separate game for each one. Thus, the total number of games a player $i$ takes part in equals their hyperdegree $k_{i}$. We define $k_{i}^{/}$ and $k_{i}^{\Delta}$ as the number of 2-player and 3-player games player $i$ is involved in, respectively, satisfying $k_{i}^{/}+k_{i}^{\Delta}=k_{i}$.
@@ II.1 Payoffs
where the first entry in each cell denotes the payoff earned by the row player, while the second entry is the payoff for the column player. $R^{/}$ denotes the reward for mutual cooperation, while $P^{/}$ denotes the penalty for mutual defection. A cooperator earns a payoff $S^{/}$ (namely sucker’s payoff) against a defector, while the defector earns a payoff $T^{/}$ (temptation to defect) when facing a cooperator. The relative order of these payoffs determines the type of 2-player social dilemmas. In particular for our analysis, we will study prisoner’s dilemma games, defined by the ordering $T^{/}>R^{/}>P^{/}>S^{/}$. To reduce the number of parameters in the game, we fix $R^{/}=1$ and $P^{/}=0$ to define the limit of the parameters as is typically done in the literature Szabó and Fáth (2007); Guo et al. (2021); Wang et al. (2023); Civilini et al. (2024); Wang et al. (2024b).
@@
A social dilemma is a type of collective action problem where there is a tension between personal benefit and collective good. For instance, a $2$-player Prisoner’s dilemma represents a social dilemma since, given the payoff ordering $T^{/}>R^{/}>P^{/}>S^{/}$ defining the game, the rational choice (or Nash equilibrium) for a player is to choose defection, even though it would be more beneficial (i.e., it would bring a higher payoff) to cooperate for both players. However, generalizing the concept of social dilemma to multi-player games is not trivial. Over the years, different definitions Kerr et al. (2004); Peña et al. (2016) of social dilemmas in multi-player game representation have been proposed. Here, we adopt the definition proposed by Ref. Broom et al. (2019), according to which the payoffs of a $3$-player game have to satisfy the following conditions to qualify as a social dilemma:
@@
However, in the higher-order game, two additional payoffs, $G^{\Delta}$ and $W^{\Delta}$, are present. If $G^{\Delta}<W^{\Delta}$, all the above conditions (a-d) hold, classifying the game as a strong social dilemma. Conversely, when $G^{\Delta}>W^{\Delta}$, condition d is not met, leading to what is known as a relaxed social dilemma. In general, the quantity $\alpha=W^{\Delta}-G^{\Delta}$ can be interpreted as a measure of the strength of the higher-order Prisoner’s Dilemma (PD). Notably, the strong social dilemma has only full defection ($\{D,D,D\}$) as a Nash equilibrium, whereas in the relaxed social dilemma, besides full defection, the strategy profile $\{C,C,D\}$ also forms a Nash equilibrium Gokhale and Traulsen (2010); Hilbe et al. (2014); Civilini et al. (2024).
@@
It is crucial that payoffs for hyperedges of different sizes remain comparable so that larger groups do not give an undue advantage (or disadvantage) simply because of their group interaction Capraro and Barcelo (2015); Peña and Nöldeke (2018); Pereda et al. (2019); Wang et al. (2024a). For instance, consider a situation where a player cooperates with two cooperating players in two dyadic interactions earning a total payoff of $2R^{/}$. If instead the player cooperated with the same two cooperating players but in a 3-player interaction, it will earn a payoff of $R^{\Delta}$. Since the evolutionary dynamics typically consists of aggregating payoffs to determine the fitness of strategies, if $2R^{/}\neq R^{\Delta}$, the total payoff earned by a player by interacting with the same players but with different types of interaction (pairwise or higher-order) will be different. Even though larger groups can have synergistic effects Capraro and Barcelo (2015); Peña and Nöldeke (2018); Pereda et al. (2019); Lee et al. (2023) making cooperation naturally more or less advantageous, we want to focus on emergent collective behaviour in absence of these phenomena. Inspired by this line of thinking, one can imagine that from a player’s perspective, a single 3-player interaction is equivalent to two pairwise interactions. Here, we introduce a new payoff constraint for the pairwise and higher-order payoffs, $R^{\Delta}\sim R^{/}+R^{/}=2R^{/}$ and similarly for $S^{\Delta}=2S^{/}$, $T^{\Delta}=2T^{/}$, and $P^{\Delta}=2P^{/}$. We tune the values of $G^{\Delta}$ and $W^{\Delta}$ to explore the landscape of higher-order PD based on its strength mentioned above.
@@ II.2 Topological overlap in higher-order networks
We consider a random-regular hypergraph such that a player $i$ has $k^{/}$ 2-hyperedges (or pairwise neighbours) and further participates in $k^{\Delta}$ 3-hyperedges (or triangles). For our simulations, we fix $k^{/}=4$ and $k^{\Delta}=2$ so that each player interacts with maximum 8 (=$k^{/}+2k^{\Delta})$ other unique players. We can represent any form of higher-order interactions as a multilayer hypergraph, where each layer consists of interactions of a specific size/order, as illustrated in fig. (1). The advantage of such a representation is that we can adapt well-established measures from the theory of multiplex networks to characterize the structure of our system. We are interested in exploring how topological similarity between the pairwise and higher-order interactions impacts the emergence of cooperation in structured populations Battiston et al. (2014); Lee et al. (2021); Anwar and Ghosh (2022); Zhang et al. (2023); Presigny et al. (2024); Malizia et al. (2025); Lamata-Otín et al. (2025).
We denote $\mathcal{E}^{\Delta}_{\text{proj}}$ as the set of projection of 3-hyperedges unto their corresponding $2$-hyperedges. In other words, if $e_{lmn}\in\mathcal{E}^{\Delta}$, then $e_{lm},e_{mn},e_{ln}\in\mathcal{E}^{\Delta}_{\text{proj}}$. We then define the topological overlap $\omega$ between $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}$ as the size of the intersection between the sets of edges $\mathcal{E}^{/}$ and $\mathcal{E}^{\Delta}_{\text{proj}}$ normalized by the size of set of edges in the pairwise network Battiston et al. (2014, 2017); Krishnagopal and Bianconi (2023); Lamata-Otín et al. (2024),
@@ II.3 Evolutionary dynamics
To account for the dynamical coupling between interaction orders, we assume that the focal player can imitate the strategies of the model player associated with the two layers. To illustrate with an example, consider the update of the pairwise strategy of a focal player $f$ denoted by $s_{f}^{/}$ (fig. (1)). We now assume that the player can copy the pairwise strategy of the model player $m$ denoted by $s_{m}^{/}$. However, here we propose that the dynamical coupling between the two layers also enables the focal player to copy the higher-order strategy of the model player denoted by $s_{m}^{\Delta}$. We denote the probability of imitating the strategy of the model player from the other layer as $p_{\text{switch}}$, controlling the dynamical coupling between the two layers and allowing for inter-order imitation. The profile of all transition probabilities at a given time step can be summarised as follows,
@@
We obtain one full Monte-Carlo Step (MCS) by repeating the above procedure $2\cdot N$ times so that each player gets the opportunity to update both the pairwise and higher-order strategies once on average. The structural and dynamical components of our model are fully visualized in fig. (1).
We characterize the pro-social behaviour in our system using several descriptors. We introduce $\rho_{0}^{/}$ and $\rho_{0}^{\Delta}$ as the initial density of cooperators in the pairwise and higher-order layers respectively. Correspondingly, $\rho_{0}=\frac{1}{2}[\rho_{0}^{/}+\rho_{0}^{\Delta}]$ denotes the initial mass of cooperators in the system. We always consider $\rho_{0}^{/}=\rho_{0}^{\Delta}=\rho_{0}=0.5$ unless stated otherwise. To quantify the stationary state properties of the system, i.e. a state where the observables of the system become time independent, we compute $\rho^{/}$ as the fraction of nodes cooperating in the pairwise interactions and $\rho^{\Delta}$ as the fraction of nodes cooperating in the higher-order interactions. We denote the overall cooperation level in the system as Wang et al. (2015); Battiston et al. (2017),
@@ III.1 Inter-order dynamical coupling mediates cooperation
We first investigate the effect of the dynamical coupling between the different orders of interactions in our model by tuning $p_{\text{switch}}$. Figure (2) depicts the stationary state cooperation levels $\langle\rho\rangle$ as a function of $p_{\text{switch}}$ for various values of social dilemma strength $\alpha$ and structural overlap $\omega$ in multi-dimensional strategy systems. To highlight the advantages of our model, we also plot in coloured dotted lines, the cooperation levels in a similar system except with a single/scalar strategy assigned to each node instead of a vector of strategies. In other words, we showcase the difference between the presence and absence of group-size-dependent strategies. Note that in the case where each node has only one associated strategy, the notion of $p_{\text{switch}}$ is meaningless. First, we observe that depending on the value of $\alpha$, the uni-dimensional scalar-strategy system shows a transition from full defection to high levels of cooperation Civilini et al. (2024). Second, we notice that for both values of overlap $\omega=0.0$ and $\omega=1.0$, for $p_{\text{switch}}=0$, the system displays 50% cooperative agents. We can understand this behaviour by noticing that the optimum strategy for the pairwise PD game based on our chosen payoff values is pure defection, while for the higher-order game, since $\alpha<0$, one of the Nash equilibria is to cooperate (see Sec. (II.1)) Gokhale and Traulsen (2010); Civilini et al. (2024).